# Universally Optimal Periodic Configurations in the Plane

### v3

@@ 1 Introduction and Overview of Results
Let $F:\mathbb{R}^{d}\rightarrow(-\infty,\infty]$ be a lower-semicontinuous potential.
@@
If for some lattice $\Lambda\subseteq\mathbb{R}^{d}$ (a lattice is a discrete, additive, rank $d$ subgroup of $\mathbb{R}^{d}$, more on lattices in Section 2.1), $F$ is $\Lambda$-periodic (i.e., $F(\cdot+v)=F$ for all $v\in\Lambda$), then we also refer to the $F$-energy as periodic energy. In this case, without loss of generality, we may assume that $\omega_{n}$ lies in the flat tori given by some specified fundamental domain $\Omega_{\Lambda}:=\mathbb{R}^{d}/\Lambda$, since replacing a point $x\in\omega_{n}$ with any point in $x+\Lambda$ does not change $E_{F}(\omega_{n})$.
@@
and for $n\not\in N$, the definition is extended by interpolation. An $n$-point configuration $\omega_{n}\subset\mathbb{R}^{d}$ satisfying $E_{F}(\omega_{n})=\mathcal{E}_{F}(n)$ is called $F$-optimal. Note that the lower-semicontinuity of $F$ and the compactness of $\Omega_{\Lambda}$ in the torus topology imply the existence of at least one $F$-optimal configuration for each cardinality $n>1$.
@@
The potential $F_{f,\Lambda}$ has the following physical interpretation: if $f(r^{2})$ represents the energy required to place a pair of unit charge particles a distance $r$ from each other, then $F_{f,\Lambda}(x)$ is the energy required to place such a particle at the point $x$ in the presence of existing particles at points of $\Lambda$. We write the pair interaction in terms of the distance squared in order to be compatible with the notion of universal optimality discussed below.
The periodization of Gaussian potentials $f_{a}(r^{2}):=\exp(-ar^{2})$ for $a>0$ leads to a type of lattice theta function (cf. [4, Chapter 10]) and plays a central role in our analysis. For convenience, we write $F_{a,\Lambda}:=F_{f_{a},\Lambda}$ or just $F_{a}$ when the choice of lattice is unambiguous.
@@
    We say $\Lambda$ is universally optimal if for any sublattice $\Phi\subseteq\Lambda$, the configuration [^1] A fundamental domain for a group $G$ acting on a set $X$ is a subset of $X$ consisting of exactly one point from each $G$-orbit. Note that $X/G$ will be used to denote both a fundamental domain and the set of $G$ orbits in $X$. In particular, if $\Omega_{\Phi}$ is a choice of fundamental domain for $\mathbb{R}^{d}/\Phi$, then $\Lambda\cap\Omega_{\Phi}$ is a choice of fundamental domain for $\Lambda/\Phi$ in (3). The index of $\Phi$ in $\Lambda$ is $[\Lambda,\Phi]:=|\Lambda/\Phi|$.
@@
If $\omega_{n}$ is $\Lambda$-universally optimal, then it follows from a theorem of Bernstein [2] (see [8],[4]) that $\omega_{n}$ is $F_{f,\Lambda}$-optimal for any $f$ with $d$-rapid decay that is completely monotone on $(0,\infty)$. [^2] Recall that a function $g$ is completely monotone on an interval $I$ if $(-1)^{n}g^{(n)}\geq 0$ on $I$ for all positive integers $n$.
@@ 2.1 Preliminaries: Lattices and Fourier Series
[…]
@@
with Fourier coefficients $\hat{g}_{v}:=\frac{1}{|\Lambda|}\int_{\Omega_{\Lambda}}g(x)$ for $v\in\Lambda^{*}$ where equality (and the implied unconditional limit on the right hand side) holds in $L^{2}(\Omega_{\Lambda})$. Of course, elements of $L^{2}(\Omega_{\Lambda})$ are actually equivalence classes of functions. If $g\in L^{2}(\Omega_{\Lambda})$ contains an element of $C(\mathbb{R}^{d})$, then we identify $g$ with its continuous representative and write $g\in L^{2}(\Omega_{\Lambda})\cap C(\mathbb{R}^{d})$. As will be the case in our applications, if $g\in L^{2}(\Omega_{\Lambda})$ is such that $\sum_{v\in\Lambda^{*}}|\hat{g}_{v}|<\infty$, then the right-hand side of (4) converges uniformly and unconditionally to $g$ and so $g\in L^{2}(\Omega_{\Lambda})\cap C(\mathbb{R}^{d})$ and (4) holds pointwise for every $x\in\mathbb{R}^{d}$.
[…]
@@ 2.2 Lattice symmetry, symmetrized basis functions, and polynomial structure
We next consider the case of a rectangular lattice by which we mean a lattice of the form $\Lambda_{R}=(a_{1}\mathbb{Z})\times\cdots(a_{d}\mathbb{Z})$ with $a_{1},\ldots,a_{d}>0$. The symmetry group of a rectangular lattice in $\mathbb{R}^{d}$ contains the subgroup $H$ of order $2^{d}$ generated by the coordinate reflections
@@
Recall the $\ell$th Chebyshev polynomial of the first kind defined by $\cos(\ell\theta)=T_{\ell}(\cos\theta)$ for $\ell=0,1,2,\ldots$. We then have the following proposition.
@@
For any $\Lambda_{R}$-periodic function $h$ with $H$-symmetry, $\tilde{h}$ will refer to the function defined on $[-1,1]^{d}$ by
@@
are polynomials in the variables $t_{1},\dots,t_{d}$. It then follows that the collection of polynomials $\{P_{v}^{\Phi}\mid v\in\Phi^{*}/G_{\Phi}\}$ is orthogonal with respect to the measure $(1-t_{1}^{2})^{-1/2}\cdots(1-t_{d}^{2})^{-1/2}dt_{1}\,\cdots\,dt_{d}$ on $[-1,1]^{d}$. Furthermore, $\tilde{h}$ is CPSD if and only if its expansion in terms of these polynomials has coefficients that are non-negative and summable.
@@ 2.3 Linear Programming Bounds for Periodic Energy
The next proposition follows immediately from (14) and the condition (15) for equality in (14). The calculations in (14) are similar to the proof of the linear programming bounds for energy found in [8, Proposition 9.3] and is closely related to Delsarte-Yudin energy bounds for spherical codes (cf. [4, Chapters 5.5 and 10.4]).
@@
and when $\Lambda=L$, we’ll consider the representative set $[0,1/2]\times[0,\sqrt{3}/2]$.
@@ 2.4 Moments for certain lattice configurations
We consider moments of configurations $\omega(\Phi,\Lambda):=\Lambda\cap\Omega_{\Phi}$ obtained by restricting a lattice $\Lambda$ to the fundamental domain of a sublattice $\Phi$ [^4] We are aware of similar lattice computations in discrete harmonic analysis (e.g., see [29]), but the authors could not find a reference for this exact result and so include a proof.
@@
Let $\Lambda=V\mathbb{Z}^{d}$; i.e., $V$ is a generator for $\Lambda$. Since $\Phi$ is a sublattice of $\Lambda$, there is some integer $d\times d$ matrix $W$ such that $VW$ is a generator for $\Phi$. Then $W$ can be written in Smith Normal Form as $W=SDT$ where $S$ and $T$ are integer matrices with determinant ±$\pm 1$ (equivalently, their inverses are also integer matrices) and $D$ is a diagonal matrix with positive integer diagonal entries $\lambda_{1},\ldots,\lambda_{d}$. It follows that $\widetilde{V}=VS$ is a generator for $\Lambda$ and $U=\widetilde{V}D$ is a generator for $\Phi$. Choosing the fundamental domains $\Omega_{\Lambda}=\widetilde{V}[0,1)^{d}$ and $\Omega_{\Phi}=U[0,1)^{d}$ gives
@@
where $[0..p]:=\{0,1,2,\ldots,p-1\}$ positive integers $p$. Let $v\in\Phi^{*}$ so that $v=U^{-T}k={\widetilde{V}}^{-T}D^{-1}k$ for some $k=(k_{1},k_{2},\dots,k_{d})\in\mathbb{Z}^{d}.$ Then $v\cdot(\widetilde{V}j)=j\cdot(D^{-1}k)$ and so
@@
=$=\begin{cases}\lambda_{1}\cdots\lambda_{d},&k\in D\mathbb{Z}^{d},\\ 0,&\text{otherwise},\end{cases}$
@@ 2.5 Lattice theta functions
and moreover, as shown below, $\tilde{\theta}$ is absolutely monotone on $[-1,1]$. First, we recall the Jacobi triple product formula.
Theorem 9.
Jacobi Triple Product Formula Let $z,q\in\mathbb{C}$ with $\lvert q\rvert<1$ and $z\neq 0$. Then
@@
For any $c>0$, the function $\tilde{\theta}=\tilde{\theta}(c;\cdot):[-1,1]\rightarrow(0,\infty)$ is strictly absolutely monotone on $[-1,1]$ and its logarithmic derivative $\tilde{\theta}^{\prime}/\tilde{\theta}$ is strictly completely monotone on $[-1,1]$.
@@
$\begin{split}F_{a,\Phi}(x)&=\sum_{v\in\Phi}e^{-a\|x+v\|^{2}}=\sum_{k\in\mathbb{Z}^{d}}\prod_{i=1}^{d}e^{-aa_{i}^{2}(x_{i}/a_{i}+k_{i})^{2}}\\ &=\prod_{i=1}^{d}\left(\sum_{k_{i}\in\mathbb{Z}}e^{-aa_{i}^{2}(\frac{x_{i}}{a_{i}}+k_{i})^{2}}\right)=\prod_{i=1}^{d}F_{aa_{i}^{2},\mathbb{Z}}(x_{i}/a_{i}).\end{split}$ (24)
@@ 2.6 Polynomial interpolation and linear programming bounds for lattice configurations
$E_{F}(\omega_{n})\geq E_{g}(\omega_{n})=n^{2}c_{0}-n\tilde{g}_{(}1,\ldots,1),$ (27)
@@
We now consider sufficient conditions for the $F$-optimality of configurations of the form $\omega(\Phi,\frac{1}{m}\Lambda)$, where $[\Lambda:\Phi]=\kappa$. For such a configuration (when the choices of $\Lambda,\Phi,$ and $\Delta_{\Phi}$ are clear), we define
@@ 2.7 Example: Universal optimality of $\mathbb{Z}$
and form the family of monic orthogonal polynomials with respect to the measure $(1-t^{2})dt$ on $[-1,1]$. These polynomials can be related to Chebyshev polynomials of the first kind through the relations
@@
Observe that the points $-1<t_{1,m}<\cdots<t_{m-1,m}<1$ are also the roots of $U_{m}$. It then follows using the Christoffel-Darboux formula that the partial products $\prod_{\ell=0}^{j}(t-t_{\ell})$ have expansions in $U_{0},U_{1},\ldots,U_{j}$ with positive coefficients for $j=1,\ldots,m-2$ (see [8, Prop 3.2] or [4, Thm A.5.9]). Likewise, $\prod_{\ell=1}^{m-1}(t-t_{\ell})$ is simply $U_{m-1}$ and $(t+1)=(t-t_{0})$, both of which are positive definite. Hence, each such partial product is PSD as is any product of such partial products; in particular, with $T=\{t_{0},t_{0},t_{1},t_{1},\ldots,t_{m-1},t_{m-1}\}$, the partial products $p_{j}(T;t)$ defined in (79) are PSD for $j\leq 2m$.
By Proposition 10, the function $\tilde{F}_{a,\mathbb{Z}}$ is absolutely monotone on $[-1,1]$ and, since the divided differences of an absolutely monotone function are non-negative, it follows that the interpolant $H_{T}(\tilde{F}_{a,\mathbb{Z}})(t)$ defined in (78) is PSD. Moreover, $H_{T}(\tilde{F}_{a,\mathbb{Z}})(t)$ has degree at most $12m-$ 1, and the error formula (81) shows that $H_{T}(\tilde{F}_{a,\mathbb{Z}})\leq\tilde{F}_{a,\mathbb{Z}}$ on [-1,1].
@@
For any nonnegative integers $l_{1}$ and $l_{2}$ whose sum $l_{1}+l_{2}$ is even, we have on all $[-1,1]^{2}$ that
@@
On all of $[-1,1]\times[\frac{1}{2},1]$, we have the inequalities
@@
As observed in [33] (also see [4, Chapter 10] and [16]), this Lemma 15 suffices to proves the $A_{2}$-universal optimality of the 2 and 3-point configurations discussed in the introduction (see section 3.5 for more detail).
@@
and $[0,1/2]\times[0,\sqrt{3}/2]$ for $\mathbb{R}^{2}/S_{L}$. Likewise for $A_{2}$, we take the sets
@@
=$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,k_{1}\equiv k_{2}\bmod 2,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\}$ (35)
@@
=$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{1}\equiv j_{2}(\bmod 2)\}$ (37)
@@
=$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,k_{1}\equiv k_{2}\bmod 2,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{2}\equiv 0\bmod 3\}$ (39)
@@
=$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{1}\equiv j_{2}(\bmod 2),j_{2}\equiv 0\bmod 3\}.$ (41)

### v4

@@ 1 Introduction and Overview of Results
Let $F:\mathbb{R}^{d}\rightarrow(-\infty,\infty]$ be a lower-semicontinuous potential. For a finite multiset $\omega_{n}=\{x_{1},...,x_{n}\}\subseteq\mathbb{R}^{d}$ of cardinality $n$, we consider the $F$-energy of $\omega_{n}$ defined by
@@
If for some lattice $\Lambda\subseteq\mathbb{R}^{d}$ (a lattice is a discrete, additive, rank $d$ subgroup of $\mathbb{R}^{d}$, more on lattices in Section 2.1), $F$ is $\Lambda$-periodic (i.e., $F(\cdot+v)=F$ for all $v\in\Lambda$), then we also refer to the $F$-energy as periodic energy. In this case, without loss of generality, we may assume that $\omega_{n}$ lies in the flat torus given by some specified fundamental domain $\Omega_{\Lambda}:=\mathbb{R}^{d}/\Lambda$, since replacing a point $x\in\omega_{n}$ with any point in $x+\Lambda$ does not change $E_{F}(\omega_{n})$.
@@
and for $n\not\in\mathbb{N}$, the definition is extended by interpolation. An $n$-point configuration $\omega_{n}\subset\mathbb{R}^{d}$ satisfying $E_{F}(\omega_{n})=\mathcal{E}_{F}(n)$ is called $F$-optimal. Note that the lower-semicontinuity of $F$ and the compactness of $\Omega_{\Lambda}$ in the torus topology imply the existence of at least one $F$-optimal configuration for each cardinality $n>1$.
@@
The potential $F_{f,\Lambda}$ has the following physical interpretation: if $f(r^{2})$ represents the energy required to place a pair of unit charge particles at distance $r$ from each other, then $F_{f,\Lambda}(x)$ is the energy required to place such a particle at the point $x$ in the presence of existing particles at points of $\Lambda$. We write the pair interaction in terms of the distance squared in order to be compatible with the notion of universal optimality discussed below. See [12] and [13] for constructions of periodic potentials from interactions $f$ without rapid decay.
The periodization of Gaussian potentials $f_{a}(r^{2}):=\exp(-ar^{2})$ for $a>0$ leads to a type of lattice theta function (cf. [3, Chapter 10]) and plays a central role in our analysis. For convenience, we write $F_{a,\Lambda}:=F_{f_{a},\Lambda}$ or just $F_{a}$ when the choice of lattice is unambiguous.
@@
    We say $\Lambda$ is universally optimal if for any sublattice $\Phi\subseteq\Lambda$, the configuration [^1] A fundamental domain for a group $G$ acting on a set $X$ is a subset of $X$ consisting of exactly one point from each $G$-orbit. Note that $X/G$ will be used to denote both a fundamental domain and the set of $G$ orbits in $X$. In particular, if $\Omega_{\Phi}$ is a choice of fundamental domain for $\mathbb{R}^{d}/\Phi$, then $\Lambda\cap\Omega_{\Phi}$ is a choice of fundamental domain for $\Lambda/\Phi$ in (3). The index of $\Phi$ in $\Lambda$ is $\Lambda,\Phi]:=|\Lambda/\Phi|$.
@@
If $\omega_{n}$ is $\Lambda$-universally optimal, then it follows from a theorem of Bernstein [2] (see [4],[3]) that $\omega_{n}$ is $F_{f,\Lambda}$-optimal for any $f$ with $d$-rapid decay that is completely monotone on $(0,\infty)$. [^2] Recall that a function $g$ is completely monotone on an interval $I$ if $(-1)^{n}g^{(n)}\geq 0$ on $I$ for all positive integers $n$.. Furthermore, taking $a\to\infty$, it follows that a universally optimal configuration must also form the centers of an optimal sphere packing.
@@ 2.1 Preliminaries: Lattices and Fourier Series
Let $\Lambda$ be a lattice in $\mathbb{R}^{d}$ with generator $V$ and fundamental domain $\Omega_{\Lambda}$. We let $L^{2}(\Omega_{\Lambda})$ denote the Hilbert space of complex-valued $\Lambda$-periodic functions on $\mathbb{R}^{d}$ with inner product $\langle f,g\rangle=\int_{\Omega_{\Lambda}}f(x)\overline{g(x)}\,dx$. Then $\{e^{2\pi iv\cdot x}\mid v\in\Lambda^{*}\}$ forms an orthogonal basis of $L^{2}(\Omega_{\Lambda})$ yielding the Fourier expansion of a function $g\in L^{2}(\Omega_{\Lambda})$:
@@
with Fourier coefficients $\hat{g}_{v}:=\frac{1}{|\Lambda|}\int_{\Omega_{\Lambda}}g(x)\,dx$ for $v\in\Lambda^{*}$ where equality (and the implied unconditional limit on the right hand side) holds in $L^{2}(\Omega_{\Lambda})$. Of course, elements of $L^{2}(\Omega_{\Lambda})$ are actually equivalence classes of functions. If $g\in L^{2}(\Omega_{\Lambda})$ contains an element of $C(\mathbb{R}^{d})$, then we identify $g$ with its continuous representative and write $g\in L^{2}(\Omega_{\Lambda})\cap C(\mathbb{R}^{d})$. As will be the case in our applications, if $g\in L^{2}(\Omega_{\Lambda})$ is such that $\sum_{v\in\Lambda^{*}}|\hat{g}_{v}|<\infty$, then the right-hand side of (4) converges uniformly and unconditionally to $g$ and so $g\in L^{2}(\Omega_{\Lambda})\cap C(\mathbb{R}^{d})$ and (4) holds pointwise for every $x\in\mathbb{R}^{d}$.
Definition 2.2.
We say that $g\in L^{2}(\Omega_{\Lambda})$ is conditionally positive semi-definite (CPSD) if the Fourier coefficients $\hat{g}_{v}\geq 0$ for all $v\in\Lambda^{*}\setminus\{0\}$ and $\sum_{v\in\Lambda^{*}}\hat{g}_{v}<\infty$; and we say that a CPSD $g$ is positive semi-definite (PSD) if $\hat{g}_{0}\geq 0$. [^3] If $g$ is PSD in the above sense, then for any configuration $\omega_{n}=(x_{1},\ldots,x_{n})$ the matrix $G=(g(x_{i}-x_{j}))$ is positive semi-definite in the sense that $v^{T}Gv\geq 0$ for any $v$ whose components sum to 0. Conversely, Bochner’s Theorem shows that any $g$ with this property is PSD in our sense.
Note that the product of two PSD functions in $L^{2}(\Omega_{\Lambda})$ is PSD.
@@ 2.2 Lattice symmetry, symmetrized basis functions, and polynomial structure
We next consider the case of a rectangular lattice by which we mean a lattice of the form $\Lambda_{R}=(a_{1}\mathbb{Z})\times\cdots(a_{d}\mathbb{Z})$ with $a_{1},\ldots,a_{d}>0$ (also referred to as an orthorhombic lattice). The symmetry group of a rectangular lattice in $\mathbb{R}^{d}$ contains the subgroup $H$ of order $2^{d}$ generated by the coordinate reflections
@@
[…]
@@
For any $\Lambda_{R}$-periodic function $h$ with $H$-symmetry, $\tilde{h}$ will refer to the function defined on $1,1]^{d}$ by
@@
are polynomials in the variables $t_{1},\dots,t_{d}$. It then follows that the collection of polynomials $\{P_{v}^{\Phi}\mid v\in\Phi^{*}/G_{\Phi}\}$ is orthogonal with respect to the measure $(1-t_{1}^{2})^{-1/2}\cdots(1-t_{d}^{2})^{-1/2}dt_{1}\,\cdots\,dt_{d}$ on $1,1]^{d}$. Furthermore, $\tilde{h}$ is CPSD if and only if its expansion in terms of these polynomials has coefficients that are non-negative and summable.
@@ 2.3 Linear Programming Bounds for Periodic Energy
The next proposition follows immediately from (14) and the condition (15) for equality in (14). The calculations in (14) are similar to the proof of the linear programming bounds for energy found in [4, Proposition 9.3] and is closely related to Delsarte-Yudin energy bounds for spherical codes (cf. [3, Chapters 5.5 and 10.4]).
@@
and when $\Lambda=L$, we’ll consider the representative set $0,1/2]\times[0,\sqrt{3}/2]$.
@@ 2.4 Moments for certain lattice configurations
We consider moments of configurations $\omega(\Phi,\Lambda):=\Lambda\cap\Omega_{\Phi}$ obtained by restricting a lattice $\Lambda$ to the fundamental domain of a sublattice $\Phi$. [^4] We are aware of similar lattice computations in discrete harmonic analysis (e.g., see [17]), but the authors could not find a reference for this exact result and so include a proof.
@@
Let $\Lambda=V\mathbb{Z}^{d}$; i.e., $V$ is a generator for $\Lambda$. Since $\Phi$ is a sublattice of $\Lambda$, there is some integer $d\times d$ matrix $W$ such that $VW$ is a generator for $\Phi$. Then $W$ can be written in Smith Normal Form as $W=SDT$ where $S$ and $T$ are integer matrices with determinant $\pm 1$ (equivalently, their inverses are also integer matrices) and $D$ is a diagonal matrix with positive integer diagonal entries $\lambda_{1},\ldots,\lambda_{d}$. It follows that $\widetilde{V}=VS$ is a generator for $\Lambda$ and $U=\widetilde{V}D$ is a generator for $\Phi$. Choosing the fundamental domains $\Omega_{\Lambda}=\widetilde{V}[0,1)^{d}$ and $\Omega_{\Phi}=U[0,1)^{d}$ gives
@@
where $0..p]:=\{0,1,2,\ldots,p-1\}$ positive integers $p$. Let $v\in\Phi^{*}$ so that $v=U^{-T}k={\widetilde{V}}^{-T}D^{-1}k$ for some $k=(k_{1},k_{2},\dots,k_{d})\in\mathbb{Z}^{d}.$ Then $v\cdot(\widetilde{V}j)=j\cdot(D^{-1}k)$ and so
@@
$=\begin{cases}\lambda_{1}\cdots\lambda_{d},&k\in D\mathbb{Z}^{d},\\ 0,&\text{otherwise},\end{cases}$
@@ 2.5 Lattice theta functions
and moreover, as shown below, $\tilde{\theta}$ is absolutely monotone on [1,1][-1,1]. First, we recall the Jacobi triple product formula.
Theorem 9 (Jacobi Triple Product Formula).
Let $z,q\in\mathbb{C}$ with $\lvert q\rvert<1$ and $z\neq 0$. Then
@@
For any $c>0$, the function $\tilde{\theta}=\tilde{\theta}(c;\cdot):[-1,1]\rightarrow(0,\infty)$ is strictly absolutely monotone on [1,1][-1,1] and its logarithmic derivative $\tilde{\theta}^{\prime}/\tilde{\theta}$ is strictly completely monotone on [1,1][-1,1].
@@
$\begin{split}F_{a,\Phi}(x)&=\sum_{v\in\Phi}e^{-a|x+v|^{2}}=\sum_{k\in\mathbb{Z}^{d}}\prod_{i=1}^{d}e^{-aa_{i}^{2}(x_{i}/a_{i}+k_{i})^{2}}\\ &=\prod_{i=1}^{d}\left(\sum_{k_{i}\in\mathbb{Z}}e^{-aa_{i}^{2}(\frac{x_{i}}{a_{i}}+k_{i})^{2}}\right)=\prod_{i=1}^{d}F_{aa_{i}^{2},\mathbb{Z}}(x_{i}/a_{i}).\end{split}$ (24)
@@ 2.6 Polynomial interpolation and linear programming bounds for lattice configurations
$E_{F}(\omega_{n})\geq E_{g}(\omega_{n})=n^{2}c_{0}-n\tilde{g}(1,\ldots,1),$ (27)
@@
We now consider sufficient conditions for the $F$-optimality of configurations of the form $\omega(\Phi,\frac{1}{m}\Lambda)$, where $\Lambda:\Phi]=\kappa$. For such a configuration (when the choices of $\Lambda,\Phi,$ and $\Delta_{\Phi}$ are clear), we define
@@ 2.7 Example: Universal optimality of $\mathbb{Z}$
and form the family of monic orthogonal polynomials with respect to the measure $(1-t^{2})dt$ on [1,1][-1,1]. These polynomials can be related to Chebyshev polynomials of the first kind through the relations
@@
Observe that the points $1<t_{1,m}<\cdots<t_{m-1,m}<1$ are also the roots of $U_{m}$. It then follows using the Christoffel-Darboux formula that the partial products $\prod_{\ell=0}^{j}(t-t_{\ell})$ have expansions in $U_{0},U_{1},\ldots,U_{j}$ with positive coefficients for $j=1,\ldots,m-2$ (see [4, Prop 3.2] or [3, Thm A.5.9]). Likewise, $\prod_{\ell=1}^{m-1}(t-t_{\ell})$ is simply $U_{m-1}$ and $(t+1)=(t-t_{0})$, both of which are positive definite. Hence, each such partial product is PSD as is any product of such partial products; in particular, with $T=\{t_{0},t_{0},t_{1},t_{1},\ldots,t_{m-1},t_{m-1}\}$, the partial products $p_{j}(T;t)$ defined in (79) are PSD for $j\leq 2m$.
By Proposition 10, the function $\tilde{F}_{a,\mathbb{Z}}$ is absolutely monotone on [1,1][-1,1] and, since the divided differences of an absolutely monotone function are non-negative, it follows that the interpolant $H_{T}(\tilde{F}_{a,\mathbb{Z}})(t)$ defined in (78) is PSD. Moreover, $H_{T}(\tilde{F}_{a,\mathbb{Z}})(t)$ has degree at most 2m12m-1, and the error formula (81) shows that $H_{T}(\tilde{F}_{a,\mathbb{Z}})\leq\tilde{F}_{a,\mathbb{Z}}$ on [-1,1].
@@
For any nonnegative integers $l_{1}$ and $l_{2}$ whose sum $l_{1}+l_{2}$ is even, we have on all $1,1]^{2}$ that
@@
On all of $1,1]\times[\frac{1}{2},1]$, we have the inequalities
@@
As observed in [22] (also see [3, Chapter 10] and [11]), this Lemma 15 suffices to proves the $A_{2}$-universal optimality of the 2 and 3-point configurations discussed in the introduction (see section 3.5 for more detail).
@@
and $0,1/2]\times[0,\sqrt{3}/2]$ for $\mathbb{R}^{2}/S_{L}$. Likewise for $A_{2}$, we take the sets
@@
$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,k_{1}\equiv k_{2}\bmod 2,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\}$ (35)
@@
$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{1}\equiv j_{2}(\bmod 2)\}$ (37)
@@
$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,k_{1}\equiv k_{2}\bmod 2,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{2}\equiv 0\bmod 3\}$ (39)
@@
$=\{[k_{1},k_{2}/\sqrt{3}]\mid k_{1},k_{2}\geq 0,[k_{1},k_{2}]\neq m[j_{1},j_{2}]\text{ for some }j_{1}\equiv j_{2}(\bmod 2),j_{2}\equiv 0\bmod 3\}.$ (41)